if the sum of two numbers is 32 and their quotient is three, what are they?
Answer by math_helper(2461)



x + y = 32, and assume y>x
a 2nd equation is
y/x = 3
and this means y=3x. Substitute this into the first equation:
x + (3x) = 32
4x = 32
x = 8 --> y = 24
--
Ans: x=8, y=24
--
Check: 8 + 24 = 32 (ok)
24/8 = 3 (ok)